OpenStudy (adamaero):

Medal: HELP? xcosy = 539.6 xsiny = 439 xcosy + xsiny = 539.6 + 493 xcosy + xsiny = 1032.6 then?

OpenStudy (anonymous):

Are you solving for \(x\)?

OpenStudy (adamaero):

x and y

OpenStudy (anonymous):

\[x\cos y+x\sin y=978.6~~\cancel{\large\implies}~~(x\cos y)^2+(x\sin y)^2=978.6^2\] In general, \(a+b=c\) does not mean \(a^2+b^2=c^2\), but rather \((a+b)^2=c^2\). Here's what I would suggest. \[\frac{x\sin y}{x\cos y}=\tan y=\frac{439}{539.6}\] Then you solve for \(y\), back-sub and solve for \(x\).

OpenStudy (adamaero):

ohhh, that was exactly how the book does it with inverse tan; thanks!